Use the ideal gas law to find T when P = 6 atm, V = 3 L, n = 5 mol, and R = 0.0821. Round answer to 1 d.p.
Answer by math_tutor2020(3817)


Ideal Gas Law formula
PV = nRT

the variables are
P = pressure
V = volume
n = number of moles
R = ideal gas constant
T = absolute temperature in Kelvin

In this case
P = 6 atmospheres
V = 3 liters
n = 5 moles
R = 0.0821 (L*atm)/(K*mol)
T = unknown Kelvin temperature

Let's solve for T.
PV = nRT
6*3 = 5*0.0821*T
T = (6*3)/(5*0.0821)
T = 43.848965
T = 43.8 Kelvin is the approximate temperature.